INAF queue observing, 2015 Nov 03-Nov 04

Observer: S. Bisogni, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Barry Rothberg
Instrument: LUCI, LBC

Summary

* 03:10 UT the humidity is 100% and the temperature is at the dew point.

Night duration (12\xB0 el twilight) 19:00-5:10 ( 02:00-12:10 UT) = 10 h 10m

Observing time: 0 hrs 0 m

Weather time loss: 10 h 10 m (100%)

technical time loss 0 m

Data Summary

At the sunset the sky is covered, humidity 86%, very strong wind.
the dome is closed
* 03:10 UT the humidity is 100% and the temperature is at the dew point.

Project_RUN Objects Filter ExpTime Notes
Calibration MODS2 (Saracco program)

INAF queue observing, 2015 Nov 04-Nov 05

Observer: S. Bisogni, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Barry Rothberg
Instrument: LUCI

Summary

00:41 UT the humidity is 96% and the dome is completely covered by snow. Tonight it snowed very hard and we are waiting for Santa to come. The staff informed us that for tonight there is not chance at all to open.

Night duration (12\xB0 el twilight) 19:00-5:10 ( 02:00-12:10 UT) = 10 h 10m

Observing time: 0 hrs 0 m

Weather time loss: 10 h 10 m (100%)

technical time loss 0 m

Data Summary

Project_RUN Objects Filter ExpTime Notes
Calibrazioni LUCI (Ragazzoni program)

INAF queue observing, 2015 Nov 05-Nov 06

Observer: S. Bisogni, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Barry Rothberg
Instrument: LUCI

Summary

22:40 UT the humidity is still high (95%), but the weather is really sunny. The telescope operator is going to check the roof to see if we can can get rid of the melted snow on it and open, but he's still not sure about the humidity situation..
00:00 still have a lot of rime ice on the weather masts, lightning rods, railings, etc. Some ice remains on the shutter seals. Falling ice has been pounding the lounge roof and high bay all day. The TO, John Little, Jenny, Barry, and Christian agree that this is an unsafe situation in which to open. Furthermore, we are steady at ~95% RH. The rear anemometer is frozen and unresponsive.
The dome is closed

Night duration (12\xB0 el twilight) 19:00-5:10 ( 02:00-12:10 UT) = 10 h 10m

Observing time: 0 hrs 0 m

Weather time loss: 10 h 10 m (100%)

technical time loss 0 m

Data Summary

Project_RUN Objects Filter ExpTime Notes
Calibration LUCI (Iovino program   HK, zJ
Bias LBC

INAF queue observing, 2015 Nov 06-Nov 07

Observer: S. Bisogni, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Dave Thompson on call
Instrument: LBC-MODS

Summary

23:51 The TO is opening the chamber very early so to ventilate and have good temperature for the beginning of the night. At the sunset the sky is clear, the humidity is only 13%

Night duration (12\xB0 el twilight) 19:00-5:10 ( 02:00-12:10 UT) = 10 h 10m

Observing time: 9 hrs 10 m (90%)

Weather time loss: 0 h (0%)

technical time loss 1 h (10%) (problem with MODS2 acquisition)

Data Summary

Project_RUN Objects Filter ExpTime Notes
standard S112 g B, r V    
Clementini PegasusIII, PiscesII B, V 0.5 ob from number 6 to 10
Bellazzini O 339.0+09.0 g,r 0.2 h completed
Bellazzini Q 342.1+20.6 g,r, 0.2 h completed
Bellazzini P 341.7+07.7 g,r 0.2h completed
Saracco M003XLSms1 red grating 2.5h  
Standard gd71 red grating 10min  
Bellazzini A 063.7+33.3 g,r 0.2h completed
Bellazzini E 092.1+09.5 g,r 0.2 h completed
Bellazzini B 084.4+24.0 g,r 0.2 h completed
Bellazzini C 086.4+31.8 g,r 0.2 h completed
Bellazzini D 090.9+30.9 g,r 0.2 h completed
Bellazzini F 100.0+36.7 g,r 0.2 h completed
Bellazzini G 100.9+09.2 g,r 0.2h during the sunset
Skyflat LBC U_spec B, V i

INAF queue observing, 2015 Nov 07-Nov 08

Observer: S. Bisogni, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Barry Rothberg on call
Instrument: LBC-MODS-LUCI

Summary

At the sunset the sky is clear, the humidity is 20%

Night duration (12\xB0 el twilight) 19:00-5:10 ( 02:00-12:10 UT) = 10 h 10m

Observing time: 8 hrs 50 m (~90%)

Weather time loss: 0 h (0%)

technical time loss 1 h 20 m (~10%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Clementini PegasusIII, PiscesII B, V 0.3 ob from number 11 to 13, average seeing 1.3-1.5'', OUT of SPECS
Cusano AndXXI B, V 15min ob 1 and 2
Cusano AndXIX B, V 15min ob 1 and 2
Cusano And_field_1 B, V 0.5h  
Antoniucci v350cep zJspec, HKspec 0.6h (spec) completed
Antoniucci v350 cep J (ima) 0.1 h completed
Antoniucci tell v350cep HIP111257 zJspec, HKspec    
Antoniucci xztau zJspec, HKspec 0.6h (spec) completed
Antoniucci xztau J (ima) 0.1 h completed
Antoniucci tell xztau HIP30393 zJspec, HKspec
Bellazzini H 104.4+04.1 g, r 0.2h completed
Castellano MACSJ0717.5 Uspec, V 36 min  
Phot Stand SA98 SVBg_rVi    
Skyflat   V, I
Antoniucci cal   zJspec, HKspec

INAF queue observing, 2015 Nov 08-Nov 09

Observer: S. Bisogni, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Barry Rothberg on call
Instrument: LBC-MODS

Summary

At the sunset the sky is clear, the humidity is 40%.

Night duration (12\xB0 el twilight) 19:00-5:20 ( 02:00-12:20 UT) = 10 h 10m

Observing time: 10 hrs 10 m (%)

Weather time loss: 0 h (%)

technical time loss 0 h (%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
SkyFLAT LUCI2   J
Standard SA112 SDT_Uspec B V g, V r i    
Clementini PegasusIII, PiscesII B, V 0.5 h ob from number 11 to 15
Saracco M003XLSms1 red grating 2.5 h completed
Saracco M004XLSms2 red grating 30 m  
Castellano MACSJ0717.5 SDT_Uspec, V_bessel 2 h 12 m up to ob 6, the last ob taken during the sunset
Dark LUCI2
-- %USERSIG{RobertaCarini - 2015-11-04}%

Comments

 
Topic revision: r7 - 08 Jan 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback